OpenStudy (anonymous):

find the product write your answer in exponential form 7^-9*7^2

OpenStudy (anonymous):

\[ 7^(-9*7^2)\]

OpenStudy (anonymous):

this is the question, right ?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

you wan the answer ?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

\[7^{-9}*7^{2}\]

OpenStudy (anonymous):

oh then it will be \[7 ^ (2-9) = 7^(-7) = \frac{ 1 }{ 7^7} = \frac{ 1 }{ 823543 }\]

OpenStudy (anonymous):

are you sure from your question ?

OpenStudy (anonymous):

i type the answer wrong sorry

OpenStudy (anonymous):

\[8^{4}*8^{-3}\]

OpenStudy (anonymous):

hhh okay

OpenStudy (anonymous):

\[8^(4−2)=8^2=64\]

OpenStudy (anonymous):

\[3^{4}*3^{8}*3^{-1}\]

OpenStudy (anonymous):

exponential form

OpenStudy (anonymous):

3^11

OpenStudy (anonymous):

\[2^{-8}*2\]

OpenStudy (anonymous):

2^-7

OpenStudy (anonymous):

\[8^{-3}*8^{2}\]

OpenStudy (anonymous):

8^-1

OpenStudy (anonymous):

\[3^{4}*3^{2}\]

OpenStudy (anonymous):

3^6

OpenStudy (anonymous):

\[2^{9}=blank*2^{6}\]

OpenStudy (anonymous):

blank = 2^3

OpenStudy (anonymous):

hey do you know how we got the results ??

OpenStudy (anonymous):

\[4^{6}*4^{3}\]

OpenStudy (anonymous):

do you know how we got the results ??

OpenStudy (anonymous):

kinda

OpenStudy (anonymous):

do you want me to explain it for you ?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

4\[4^{6}*4^{2}\]

OpenStudy (anonymous):

sure look, this thing is very simple and very important you have only TWO rules 1- if the operation was MULTIPLICATION then we ADD the powers of them 2- if the operation was DIVISION then we SUBTRACT the powers of them

OpenStudy (anonymous):

so can you find out (\[4^6∗4^2\]) ?

OpenStudy (anonymous):

no we add the POWERS only

OpenStudy (anonymous):

\[9^{2}\]

OpenStudy (anonymous):

no the (base)(4) stay without changing only the powers are added so it will be 4^8

OpenStudy (anonymous):

we add 2+6 to get 8 and the 4 stay without changing

OpenStudy (anonymous):

\[8^{-2}*8^{-9}\]

OpenStudy (anonymous):

you do it

OpenStudy (anonymous):

and ill check it

OpenStudy (anonymous):

do we add the negative powers

OpenStudy (anonymous):

is it multiplication or division ?

OpenStudy (anonymous):

yea multiplication

OpenStudy (anonymous):

then we add the powers

OpenStudy (anonymous):

\[11^{8}\]

OpenStudy (anonymous):

is that the answer

OpenStudy (anonymous):

yes, just flip it to be : 8 ^11

OpenStudy (anonymous):

that was wrong

OpenStudy (anonymous):

no you got the idea

OpenStudy (mathmale):

@kylehenderson: Why not read through this conversation and then attempt the problem on y our own? That will determine whether you've learned this material or not.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

select the set that is not a subset of x= a,f,h,m,s